In the mercury spectrum one can find a strong emission line with the wavelength
576,960 nm. This corresponds to the energy difference between two energy levels in the mercury atom. Determine this energy difference expressed in electron volts.
1
Expert's answer
2020-05-15T08:43:38-0400

According to the Planck's formula, the energy correspondes to the wavelength "\lambda = 576,960 nm" (that is, the energy difference between two energy levels in the mercury atom) will be:

"E = h\cdot \dfrac{c}{\lambda}",

where "h = 4.14\cdot 10^{-15}\space eV\cdot s" is the Planck's constant and "c = 3\cdot 10^8 m/s" is the speed of light. Thus:

"E = 4.14\cdot 10^{-15}\space \cdot \dfrac{3\cdot 10^8}{576.960 \cdot 10^{-9}} \approx 2.15 eV" .


Answer. E = 2.15 eV.